OpenStudy (anonymous):

what is the factored form of the expression? k^2-81h^2 please help!!!

OpenStudy (reemii):

\(a^2-b^2 = (...)\times (...)\) It is a pretty common formula. \(a\) would be your \(k\) and \(b\) would be ... ?

OpenStudy (anonymous):

i have these posible answers : a.h^2(k+9)(k-9) b.(k9h^2)(k+9) c.(k+9h)(k+9h) d.(k+9h)(k-9h)

OpenStudy (reemii):

then, 2 options: check the four answers they gave you, or remember the formula and find the correct answer afterwards. The formula is the factorization of the "difference of two squares". You should try remembering it. \(b^2=811h^2\) therefore \(k^2-81h^2=(k+9h)(k-9h)\). Compare withe the formula using \(a\) and \(b\).

OpenStudy (anonymous):

got it ...thank you very much (: